# GUTs - how common are they?

### v1

@@ III Discussion
To check whether the SM gauge algebra itself is special when it comes to unifiability, we repeated the analysis for different semi-simple parts of the gauge algebra. In the case of $SU(3)\times U(1)$, we need to go to $D\geq 27$ to find the first non-VL, unifiable representation, where we find a unifiable fraction of order $10^{-5}$, similar to the SM result, even if we allow for $\tilde{S}=7$ in order not to exclude any unifiable representations (see the right panel of Fig. 2). For the remaining rank-2 gauge algebras $\{SO(5),SU(2)\times SU(2),SP(4),G_{2}\}\times U(1)$ we do not find any chiral, unifiable fermion representations with $D\leq 30$. The same is true for rank-3 algebras, such as $SU(4)\times U(1)$, with the exception of the SM gauge algebra. This is partially due to the increasing dimension of the smallest representations of these algebras. At the same time the base set of chiral theories is still growing exponentially since $U(1)$ charge assignments can be used to make sets of five or more fermions chiral [12, 13]. Thus, the number of chiral, unifiable representations is typically smaller than for the SM gauge algebra. Any result for the SM gauge algebra can therefore be considered conservative, since allowing for different gauge algebras would reduce the fraction of chiral, unifiable theories.
Let us also mention that different definitions of unifiability are possible. Our approach of demanding that the fermions fit neatly into a representation of the GUT that includes no further fermions may be relaxed. From a practical perspective, there is no harm if a GUT predicts hitherto unobserved VL fermions – those may be heavy. This is the case for instance for the right handed neutrinos predicted in $SO(10)$ grand unification [7]. A new condition to provide closure to the problem would be needed (eg. an arbitrary cut on the number of inferred VL fermions). Using our existing results, we can estimate the impact of relaxing the unifiability criterion by considering all theories as unifiable that unify when adding VL fermions up to a total fermion dimension of $30$: In this case, the unifiable fraction of Fig. 1 increases by a factor of up to (2) 4 in the (fully) chiral case, and the first chiral unifiable theory appears at $D=12$. We leave the study of stronger unification criteria (eg. unification of fermions into an irreducible representation, as in $SO(10)$ GUTs) to future work.
It is possible that the unifiability of the SM is not a consequence of originating from a grand unified theory but a consequence of another property that correlates with unifiability. As seen in Fig. 1, being fully chiral is one such property that is highly correlated with unifiability in the immediate neighborhood of the SM. Ironically, when adding our top-down expectations that only the fully chiral fermions stay light, the unifiability of the observed fermions is no longer by itself a strong indication for grand unification.
@@ IV Conclusion
In this letter, we consider the single-generation SM as one among many similar consistent theories. From this starting point, the observation that the SM fermion representation is unifiable may seem surprising. Here we try to find a quantitative answer to the question of how surprised we should actually be. We find that the unifiability of fermions into a representation of a simple unified algebra, as it occurs in the SM into a $SU(5)$ GUT, is rare among SM-like chiral theories at the 1/100 level. Stronger statements are possible when allowing for larger fermion dimension, charges or number of distinct irreducible representations than the SM. On the other hand, when restricting the analysis to only include fully chiral theories, unifiability is common among anomaly-free fermion representations not larger than the SM. The argument presented here can be taken as a purely group-theoretical indication for Grand Unification, conceptually similar to a naturalness argument. However, the absence of a probability measure in the space of theories hampers a probabilistic interpretation of our results.

### v2

@@ III Discussion
We now return to the question we set out to answer, namely whether the fact that the observed fermions of the SM can be unified is in itself surprising and may be a hint to what lies beyond the SM, or just a quirk of group theory common among representations of reductive Lie algebras. Different definitions of the base set of SM-like theories result in different values of the unifiable fraction, with a clear pattern. Many of the completely chiral representations of the SM gauge algebra (ie. strictly SM-like theories) with dimension and charges not much larger than the single generation SM ($D_{\mathrm{max}}\lesssim 17,\;|Q|\leq 10,\;\tilde{S}\leq 4$) can be exactly embedded into a representation of $SU(5)$ or $SO(10)$, resulting in an $\mathcal{O}(1)$ unifiable fraction. However, once we look at any more general set of consistent fermion representations beyond this immediate neighborhood of the single generation SM, unifiability becomes a rare property: When allowing the dimension of the representation to be only a little larger than that of the SM, unifiability quickly becomes rarer than $10^{-2}$ (see Fig. 1). Including partially chiral representations also pushes the unifiable fraction below $10^{-2}$ (see Fig. 3). Considering all small reductive gauge algebras with an abelian part (ie. loosely SM-like theories), only $10^{-3}$ of the minimal neighborhood (ie. representations with $D_{\mathrm{max}}\leq 15,\;|Q|\leq 6,\;\tilde{S}\leq 4$) of the single-generation SM is unifiable (see eqn. 6). This demonstrates that fermion unifiability as it exists generation-by-generation in the Standard Model is not a common property among similar quantum field theories.
Let us also mention that different definitions of unifiability are possible. Our approach of demanding that the fermions fit neatly into a representation of the GUT that includes no further fermions may be relaxed. From a practical perspective, there is no harm if a GUT predicts hitherto unobserved VL fermions – those may be heavy. This is the case for instance for the right handed neutrinos predicted in $SO(10)$ grand unification [4]. A new condition to provide closure to the problem would be needed (eg. an arbitrary cut on the number of inferred VL fermions). Using our existing results, we can estimate the impact of relaxing the unifiability criterion by considering all theories as unifiable that unify when adding VL fermions up to a total fermion dimension of $30$: In this case, the unifiable fraction of SM representations in Fig. 1 increases by a factor of up to 2, and the first partially chiral unifiable theory appears at $D=12$. We leave the study of stronger unification criteria (eg. unification of fermions into an irreducible representation, as in $SO(10)$ GUTs) to future work.
@@ IV Conclusion
In this work, we consider the single-generation SM as one among many similar consistent theories. From this starting point, the observation that the SM fermion representation is unifiable may seem surprising. Here we try to find a quantitative answer to the question of how surprised we should actually be. We find that the unifiability of fermions into a representation of a simple unified algebra, as it occurs in the SM into a $SU(5)$ GUT, is rare ($10^{-2}$) among SM-like chiral theories once we go beyond the immediate neighborhood of the single-generation SM. The argument presented here can be taken as a purely group-theoretical indication for Grand Unification, conceptually similar to a naturalness argument. However, the absence of a probability measure in the space of theories hampers a probabilistic interpretation of our results.